# ReAct Meets ActRe: Autonomous Annotations of Agent Trajectories for Contrastive Self-Training

### v1

@@ Abstract
Language agents have demonstrated autonomous decision-making abilities by reasoning with foundation models. Recently, efforts have been made to train language agents for performance improvement, with multi-step reasoning and action trajectories as the training data. However, collecting such trajectories still requires considerable human effort, by either artificial annotations or implementations of diverse prompting frameworks. In this work, we propose AT, a framework that enables the Autonomous Annotation of Agent Trajectories in the style of ReAct. The central role is an ActRe prompting agent, which explains the reason for an arbitrary action. When randomly sampling an external action, the ReAct-style agent could query the ActRe agent with the action to obtain its textual rationales. Novel trajectories are then synthesized by prepending the posterior reasoning from ActRe to the sampled action. In this way, the ReAct-style agent executes multiple trajectories for the failed tasks, and selects the successful ones to supplement its failed trajectory for contrastive self-training. Realized by policy gradient methods with binarized rewards, the contrastive self-training with accumulated trajectories facilitates a closed loop for multiple rounds of language agent self-improvement. We conduct experiments using QLoRA fine-tuning with the open-sourced Mistral-7B-Instruct-v0.2. In AlfWorld, the agent trained with AT obtains a 1-shot success rate of 96%, and 100% success with 4 iterative rounds. In WebShop, the 1-shot performance of the AT agent matches human average, and 4 rounds of iterative refinement lead to the performance approaching human experts. AT agents significantly outperform existing techniques, including prompting with GPT-4, advanced agent frameworks, and fully fine-tuned LLMs.
@@ 1 Introduction
In this work, we propose AT, a framework that enables Autonomous Annotation of Agent Trajectories in the style of ReAct (Yao et al., 2023) for self-improvement with minimal human supervision. The central idea is to exploit both the in-context language ability and the decision-making ability of a language agent: To collect diverse trajectories, an agent could randomly sample external actions from the action space at arbitrary steps. However, the corresponding reason for the sampled action should be annotated for a ReAct-style agent. To facilitate this, we propose ActRe, an act-then-reason prompting agent that explains the reason for the sampled action. With ActRe, the ReAct-style agent composes extra reason-then-act trajectories for each failed task by inversely prepending the ActRe-prompted reason to the randomly sampled action. After the execution of each composed trajectory, the agent receives a terminal reward from the environment, which automatically annotates the quality of the trajectory. The gathered successful trajectories are then supplemented with the failed trajectory by the ReAct-style agent alone for contrastive self-training, where we use policy gradient methods (Williams, 1992) with binarized rewards for LLM fine-tuning. As new agents are trained, more trajectories can be gathered and accumulated, which forms a closed loop for the self-improvement of language agents as shown in Fig. 1-(c).
We validate our AT framework in the textual embodied environment AlfWorld (Shridhar et al., 2021) and the online shopping environment WebShop (Yao et al., 2022). We use QLoRA (Dettmers et al., 2023) to fine-tune Mistral-7B-Instruct-v0.2 (Jiang et al., 2023) in the training experiments. Experimental performances demonstrate significant improvement over state-of-the-art agent techniques: On AlfWorld, our trained agent achieves a 96% success rate in unseen scenarios with a single trial. On WebShop, the success rate of our agent reaches 49%, which matches the average human performance (50%). In the setting of iterative refinement, after four rounds of data collection and contrastive self-training, the accumulated success rate becomes 100% on AlfWorld and 54.8% on WebShop, narrowing the gap with human experts (59.6% on WebShop). AT paves the way for agents with improved autonomy through the closed loop of self-annotation and contrastive self-training.
2 AT for Closed-Loop Self-Improvement
In this section, we introduce the closed-loop self-improvement for agents facilitated by the AT framework. The loop contains two parts: autonomous trajectory annotation with the ActRe agent (Sec. 2.1), and the contrastive self-training process (Sec. 2.2).
@@ 2.1 Autonomous Trajectory Annotation with ActRe
To automate the process, we propose a complementary ActRe prompting agent to synthesize the modified reasons by leveraging the in-context language ability of an LLM. ActRe inverts the causality of ReAct: while ReAct conditions the external action with a reason a priori, ActRe explains the reason a posteriori for an external action. The synergy of ActRe with ReAct facilitates autonomous annotations of textual reasons: when the language agent randomly samples an external action, the reason for the sampled action is obtained by querying the ActRe prompting agent. The synthetic reason is then used as the condition of the sampled action for the ReAct-style agent. The progress of a trajectory is synchronized between the ReAct-style agent and the ActRe prompting agent, with the only difference in the order of intermediate reasoning and actions. The detailed workings are depicted below:
@@ 2.2 Contrastive Self-Training
In implementation, we binarize the reward of the failed trajectories with . To address Remark 3, we let the agent collect multiple successful trajectories via diverse exploration to satisfy . After training, the new agent would follow Sec. 2.1 to gather more annotated trajectories. The trajectory set then continually grows as looping more rounds of data collection and agent training. For the training in each round, we use the accumulated trajectory set to fine-tune an LLM with Eq. (1). Another implementation detail is that in the initial round, we use 1-shot ReAct prompting to gather the training trajectories instead of exploration and annotation for bootstrapping. The failed trajectory for each task is directly excluded as it is not paired with sampled successful trajectories. Eq. (1) is therefore reduced to ReAct supervised fine-tuning in Yao et al. (2023) for the training in Round 0. The latter rounds leverage explored trajectories with autonomous annotations, and self-training by Eq. (1) with binarized rewards. Other details are covered in Appendix A.
@@ 3 Experiments
We conduct experiments on two benchmarks to valid the effectiveness of AT: the textual embodied environment AlfWorld (Shridhar et al., 2021), and the online shopping environment WebShop (Yao et al., 2022). The two benchmarks require a language agent to perform multi-step decision-making to accomplish a certain goal introduced in each task.
In AT, we loop for 4 rounds of trajectory collection and agent training, with the initial round using ReAct prompting as the bootstrap of training data. No trajectories are gathered from testing tasks for training. We use gpt-3.5-turbo-instruct-0914 to implement the initial ReAct prompting, as well as the ActRe prompting agent that helps the trajectory composition in the latter rounds. We use the open-sourced Mistral-7B-Instruct-v0.2 (Jiang et al., 2023) with QLoRA (Dettmers et al., 2023) finetuning for the training experiments.
We compare our AT framework with multiple strong baselines, including methods like advanced prompting frameworks using GPT-4, specialized LLMs by full fine-tuning, and gpt-3.5-turbo-1106 fine-tuning. The results are reported in the following sections.
@@ 3.1 AlfWorld
| BUTLER (Shridhar et al., 2021) | 46 | 39 | 74 | 100 | 22 | 24 | 37 |
| AgentLM (Zeng et al., 2023) | - | - | - | - | - | - | 86 |
@@
| ReAct (Yao et al., 2023) | 92 | 58 | 96 | 86 | 78 | 41 | 71 |
| ReAct (Our rerun) | 66 | 87 | 86 | 100 | 88 | 64 | 83 |
| AT (Round=0) | 96 | 77 | 100 | 95 | 94 | 47 | 86 |
| AT (Round=1) | 96 | 94 | 96 | 95 | 89 | 94 | 94 |
| AT (Round=2) | 100 | 100 | 100 | 95 | 89 | 82 | 96 |
| AT (Round=3) | 100 | 100 | 91 | 100 | 100 | 71 | 95 |
@@
| AT (Round=0) | 1 | 96 | 77 | 100 | 95 | 94 | 47 | 86 |
| AT (Round=1, accum.) | 2 | 96 | 94 | 100 | 95 | 100 | 100 | 97 |
| AT (Round=2, accum.) | 3 | 100 | 100 | 100 | 95 | 100 | 100 | 99 |
| AT (Round=3, accum.) | 4 | 100 | 100 | 100 | 100 | 100 | 100 | 100 |
@@ 3.2 WebShop
| AgentLM (Zeng et al., 2023) | - | 70.8/ - | |
| AT (Round=0) | 70.1/41.0 | 72.4/45.4 | |
| AT (Round=1) | 69.7/43.0 | 73.1/49.0 | |
| AT (Round=2) | 69.0/43.8 | 73.0/48.0 | |
| AT (Round=3) | 69.1/42.8 | 73.9/49.0 | |
| Iterative | LATS (Zhou et al., 2023a) | - | 75.9/38.0 |
| AT (Round=1, accum.) | 74.0/47.3 | 76.6/51.6 | |
| AT (Round=2, accum.) | 75.1/49.5 | 77.8/53.4 | |
| AT (Round=3, accum.) | 75.9/51.3 | 78.2/54.8 | |
@@
| Prompting | GPT-3.5-Turbo-16k-0613 | 52.8 | 50.6 | 52.2 |
| GPT-4-Turbo-0613 | 67.6 | 67.4 | 67.6 | |
| GPT-4-32k-0613 | 67.5 | 69.6 | 68.1 | |
| Training | xLAM-v0.1 (Zhang et al., 2024) | 53.2 | 51.2 | 52.4 |
| AT (Round=0) | 74.1 | 66.5 | 72.0 | |
| AT (Round=1) | 73.6 | 73.2 | 73.5 | |
| AT (Round=2) | 71.6 | 74.2 | 72.3 | |
| AT (Round=3) | 72.5 | 73.9 | 72.9 |
@@
Tables 4 and 5 demonstrate the significance of AT agents. With a single test trial, the AT agent matches averaged human performance (reward: 73.9 v.s. 75.5; success rate: 49.0% v.s. 50.0%). With 4 shots of test trials, AT achieves a 54.8% success rate, closing the gap with human expert performance (59.6%). The 1-shot AT agent also outperforms prompting with GPT-4-32k-0613 in both the easy and the hard split of WebShop from AgentBoard. Table 6 further shows the quality improvement of the accumulated trajectories across multiple rounds of AT. Case studies for annotated trajectories, as well as the dataset statistics for each round of training are reported in Appendices B and C.2, respectively.
@@
| AT (Round=0) | Supervised with | 70.1/41.0 | 72.4/45.4 |
@@
| AT (Round=1) | PG with binarized rewards | 69.7/43.0 | 73.1/49.0 |
@@ 4.1 Variants of the Self-Training Techniques
We conduct ablated experiments on WebShop, as it provides a real-valued reward ranging from 0 to 1. We first study the effect of different reward thresholds for training trajectory filtering. In AT, only the trajectories with reward are used in training. We change the constraint into and and compare the performance of the initial round with supervised ReAct fine-tuning. According to Table 7, the best performance is still obtained with , which agrees with the findings of Zhou et al. (2023b).
We proceed to ablate the policy gradient technique with binarized rewards and conduct experiments for training in Round 1. The first type to be compared with is supervised fine-tuning. We implement supervised training with the successful trajectories only (namely in Eq. (2)). This echoes the practice adopted by Singh et al. (2023). We also prepend the trajectories with the label conditions “Success” / “Fail” in the training data and conduct Round-1 supervised training. For the comparisons with policy gradient Eq. (2), we alternatively set to be the original reward provided by WebShop, or to be following the practice of Wang et al. (2024). When using the original reward, we also include another setting by relaxing the reward threshold constraint to be . The comparisons are shown in Table 7. Conclusions can be drawn that: (1) Policy gradient methods lead to higher promotion in task performance than supervised training methods. This resembles the effectiveness of RLHF (Ouyang et al., 2022) on top of supervised fine-tuning. (2) The binarized rewards () used in AT lead to a significant improvement of success rate. We leave the incorporation of advanced RL (Zhou et al., 2024) and RLAIF (Lee et al., 2023; Chen et al., 2024; Hosseini et al., 2024; Yuan et al., 2024) algorithms into AT as future work.
@@
While all of the experiments we previously reported are conducted with Mistral-7B-Instruct-v0.2 and QLoRA finetuning, in this section, we also validate AT with gpt-3.5-turbo-1106 finetuning, the proprietary service provided by OpenAI. As the initial trajectory set for Round-0 training in AT is obtained by ReAct prompting with gpt-3.5-turbo-instruct-0914, the starting point for the two base LLMs is the same.
Tables 8 and 9 report the performance comparison of Round-0 supervised training between the open-sourced and the proprietary LLMs. In AlfWorld, the performance of the QLoRA fine-tuned Mistral-7B-Instruct-v0.2 even surpasses that of the proprietary gpt-3.5-turbo-1106 fine-tuning service. In WebShop, the proprietary gpt-3.5-turbo-1106 finetuning performs better in Round-0 supervised training. We then let the two models separately compose diverse trajectories for their self-training. Because of the expense of inferring with the finetuned gpt-3.5-turbo-1106 model333The pricing is listed in https://openai.com/pricing, we compose trajectories for each failed training task ( with the open-sourced LLM). Shown in Table 10, the quality of the accumulated trajectories composed by the proprietary LLM is on par with those composed by the open-sourced LLM. After Round-1 self-training, the open-sourced model achieves an even higher test success rate. This is attributed to the proprietary service providing only the supervised fine-tuning option, while also indicating the importance of contrastive fine-tuning in AT.

### v2

# ReAct Meets ActRe: Autonomous Annotation of Agent Trajectories for Contrastive Self-Training

@@ Abstract
Language agents have demonstrated autonomous decision-making abilities by reasoning with foundation models. Recently, efforts have been made to train language agents for performance improvement, with multi-step reasoning and action trajectories as the training data. However, collecting such trajectories still requires considerable human effort, by either artificial annotation or implementations of diverse prompting frameworks. In this work, we propose A3T, a framework that enables the Autonomous Annotation of Agent Trajectories in the style of ReAct. The central role is an ActRe prompting agent, which explains the reason for an arbitrary action. When randomly sampling an external action, the ReAct-style agent could query the ActRe agent with the action to obtain its textual rationales. Novel trajectories are then synthesized by prepending the posterior reasoning from ActRe to the sampled action. In this way, the ReAct-style agent executes multiple trajectories for the failed tasks, and selects the successful ones to supplement its failed trajectory for contrastive self-training. Realized by policy gradient methods with binarized rewards, the contrastive self-training with accumulated trajectories facilitates a closed loop for multiple rounds of language agent self-improvement. We conduct experiments using QLoRA fine-tuning with the open-sourced Mistral-7B-Instruct-v0.2. In AlfWorld, the agent trained with A3T obtains a 1-shot success rate of 96%, and 100% success with 4 iterative rounds. In WebShop, the 1-shot performance of the A3T agent matches human average, and 4 rounds of iterative refinement lead to the performance approaching human experts. A3T agents significantly outperform existing techniques, including prompting with GPT-4, advanced agent frameworks, and fully fine-tuned LLMs.
@@ 1 Introduction
In this work, we propose A3T, a framework that enables Autonomous Annotation of Agent Trajectories in the style of ReAct (Yao et al., 2023) for self-improvement with minimal human supervision. The central idea is to exploit both the in-context language ability and the decision-making ability of a language agent: To collect diverse trajectories, an agent could randomly sample external actions from the action space at arbitrary steps. However, the corresponding reason for the sampled action should be annotated for a ReAct-style agent. To facilitate this, we propose ActRe, an act-then-reason prompting agent that explains the reason for the sampled action. With ActRe, the ReAct-style agent composes extra reason-then-act trajectories for each failed task by inversely prepending the ActRe-prompted reason to the randomly sampled action. After the execution of each composed trajectory, the agent receives a terminal reward from the environment, which automatically annotates the quality of the trajectory. The gathered successful trajectories are then supplemented with the failed trajectory by the ReAct-style agent alone for contrastive self-training, where we use policy gradient methods (Williams, 1992) with binarized rewards for LLM fine-tuning. As new agents are trained, more trajectories can be gathered and accumulated, which forms a closed loop for the self-improvement of language agents as shown in Fig. 1-(c).
We validate our A3T framework in the textual embodied environment AlfWorld (Shridhar et al., 2021) and the online shopping environment WebShop (Yao et al., 2022). We use QLoRA (Dettmers et al., 2023) to fine-tune Mistral-7B-Instruct-v0.2 (Jiang et al., 2023) in the training experiments. Experimental performances demonstrate significant improvement over state-of-the-art agent techniques: On AlfWorld, our trained agent achieves a 96% success rate in unseen scenarios with a single trial. On WebShop, the success rate of our agent reaches 49%, which matches the average human performance (50%). In the setting of iterative refinement, after four rounds of data collection and contrastive self-training, the accumulated success rate becomes 100% on AlfWorld and 54.8% on WebShop, narrowing the gap with human experts (59.6% on WebShop). A3T paves the way for agents with improved autonomy through the closed loop of self-annotation and contrastive self-training.
2 A3T for Closed-Loop Self-Improvement
In this section, we introduce the closed-loop self-improvement for agents facilitated by the A3T framework. The loop contains two parts: autonomous trajectory annotation with the ActRe agent (Sec. 2.1), and the contrastive self-training process (Sec. 2.2).
@@ 2.1 Autonomous Trajectory Annotation with ActRe
To automate the process, we propose a complementary ActRe prompting agent to synthesize the modified reasons by leveraging the in-context language ability of an LLM. ActRe inverts the causality of ReAct: while ReAct conditions the external action with a reason a priori, ActRe explains the reason a posteriori for an external action. The synergy of ActRe with ReAct facilitates the autonomous annotation of textual reasons: when the language agent randomly samples an external action, the reason for the sampled action is obtained by querying the ActRe prompting agent. The synthetic reason is then used as the condition of the sampled action for the ReAct-style agent. The progress of a trajectory is synchronized between the ReAct-style agent and the ActRe prompting agent, with the only difference in the order of intermediate reasoning and actions. The detailed workings are depicted below:
@@ 2.2 Contrastive Self-Training
In implementation, we binarize the reward of the failed trajectories with . To address Remark 3, we let the agent collect multiple successful trajectories via diverse exploration to satisfy . After training, the new agent would follow Sec. 2.1 to gather more annotated trajectories. The trajectory set then continually grows as looping more rounds of data collection and agent training. For the training in each round, we use the accumulated trajectory set to fine-tune an LLM with Eq. (1). Another implementation detail is that in the initial round, we use 1-shot ReAct prompting to gather the training trajectories instead of exploration and annotation for bootstrapping. The failed trajectory for each task is directly excluded as it is not paired with sampled successful trajectories. Eq. (1) is therefore reduced to ReAct supervised fine-tuning in Yao et al. (2023) for the training in Round 0. The latter rounds leverage explored trajectories via autonomous annotation, and self-training by Eq. (1) with binarized rewards. Other details are covered in Appendix A.
@@ 3 Experiments
We conduct experiments on two benchmarks to valid the effectiveness of A3T: the textual embodied environment AlfWorld (Shridhar et al., 2021), and the online shopping environment WebShop (Yao et al., 2022). The two benchmarks require a language agent to perform multi-step decision-making to accomplish a certain goal introduced in each task.
In A3T, we loop for 4 rounds of trajectory collection and agent training, with the initial round using ReAct prompting as the bootstrap of training data. No trajectories are gathered from testing tasks for training. We use gpt-3.5-turbo-instruct-0914 to implement the initial ReAct prompting, as well as the ActRe prompting agent that helps the trajectory composition in the latter rounds. We use the open-sourced Mistral-7B-Instruct-v0.2 (Jiang et al., 2023) with QLoRA (Dettmers et al., 2023) finetuning for the training experiments.
We compare our A3T framework with multiple strong baselines, including methods like advanced prompting frameworks using GPT-4, specialized LLMs by full fine-tuning, and gpt-3.5-turbo-1106 fine-tuning. The results are reported in the following sections.
@@ 3.1 AlfWorld
| BUTLER 8 (Shridhar et al., 2021) | 46 | 39 | 74 | 100 | 22 | 24 | 37 |
| AgentLM∗ (Zeng et al., 2023) | - | - | - | - | - | - | 86 |
@@
| ReAct 6 (Yao et al., 2023) | 92 | 58 | 96 | 86 | 78 | 41 | 71 |
| ReAct 6 (Our rerun) | 66 | 87 | 86 | 100 | 88 | 64 | 83 |
| A3T (Round=0) | 96 | 77 | 100 | 95 | 94 | 47 | 86 |
| A3T (Round=1) | 96 | 94 | 96 | 95 | 89 | 94 | 94 |
| A3T (Round=2) | 100 | 100 | 100 | 95 | 89 | 82 | 96 |
| A3T (Round=3) | 100 | 100 | 91 | 100 | 100 | 71 | 95 |
@@
| A3T (Round=0) | 1 | 96 | 77 | 100 | 95 | 94 | 47 | 86 |
| A3T (Round=1, accum.) | 2 | 96 | 94 | 100 | 95 | 100 | 100 | 97 |
| A3T (Round=2, accum.) | 3 | 100 | 100 | 100 | 95 | 100 | 100 | 99 |
| A3T (Round=3, accum.) | 4 | 100 | 100 | 100 | 100 | 100 | 100 | 100 |
@@ 3.2 WebShop
| AgentLM∗ (Zeng et al., 2023) | - | 70.8/ - | |
| A3T (Round=0) | 70.1/41.0 | 72.4/45.4 | |
| A3T (Round=1) | 69.7/43.0 | 73.1/49.0 | |
| A3T (Round=2) | 69.0/43.8 | 73.0/48.0 | |
| A3T (Round=3) | 69.1/42.8 | 73.9/49.0 | |
| Iterative | LATS 30 (Zhou et al., 2023a) | - | 75.9/38.0 |
| A3T 2 (Round=1, accum.) | 74.0/47.3 | 76.6/51.6 | |
| A3T 3 (Round=2, accum.) | 75.1/49.5 | 77.8/53.4 | |
| A3T 4 (Round=3, accum.) | 75.9/51.3 | 78.2/54.8 | |
@@
| Prompting | GPT-3.5-Turbo-16k-0613∗ | 52.8 | 50.6 | 52.2 |
| GPT-4-Turbo-0613∗ | 67.6 | 67.4 | 67.6 | |
| GPT-4-32k-0613∗ | 67.5 | 69.6 | 68.1 | |
| Training | xLAM-v0.1∗ (Zhang et al., 2024) | 53.2 | 51.2 | 52.4 |
| A3T (Round=0) | 74.1 | 66.5 | 72.0 | |
| A3T (Round=1) | 73.6 | 73.2 | 73.5 | |
| A3T (Round=2) | 71.6 | 74.2 | 72.3 | |
| A3T (Round=3) | 72.5 | 73.9 | 72.9 |
@@
Tables 4 and 5 demonstrate the significance of A3T agents. With a single test trial, the A3T agent matches averaged human performance (reward: 73.9 v.s. 75.5; success rate: 49.0% v.s. 50.0%). With 4 shots of test trials, A3T achieves a 54.8% success rate, closing the gap with human expert performance (59.6%). The 1-shot A3T agent also outperforms prompting with GPT-4-32k-0613 in both the easy and the hard split of WebShop from AgentBoard. Table 6 further shows the quality improvement of the accumulated trajectories across multiple rounds of A3T. Case studies for annotated trajectories, as well as the dataset statistics for each round of training are reported in Appendices B and C.2, respectively.
@@
| A3T (Round=0) | Supervised with | 70.1/41.0 | 72.4/45.4 |
@@
| A3T (Round=1) | PG with binarized rewards | 69.7/43.0 | 73.1/49.0 |
@@ 4.1 Variants of the Self-Training Techniques
We conduct ablated experiments on WebShop, as it provides a real-valued reward ranging from 0 to 1. We first study the effect of different reward thresholds for training trajectory filtering. In A3T, only the trajectories with reward are used in training. We change the constraint into and and compare the performance of the initial round with supervised ReAct fine-tuning. According to Table 7, the best performance is still obtained with , which agrees with the findings of Zhou et al. (2023b).
We proceed to ablate the policy gradient technique with binarized rewards and conduct experiments for training in Round 1. The first type to be compared with is supervised fine-tuning. We implement supervised training with the successful trajectories only (namely in Eq. (2)). This echoes the practice adopted by Singh et al. (2023). We also prepend the trajectories with the label conditions “Success” / “Fail” in the training data and conduct Round-1 supervised training. For the comparisons with policy gradient Eq. (2), we alternatively set to be the original reward provided by WebShop, or to be following the practice of Wang et al. (2024). When using the original reward, we also include another setting by relaxing the reward threshold constraint to be . The comparisons are shown in Table 7. Conclusions can be drawn that: (1) Policy gradient methods lead to higher promotion in task performance than supervised training methods. This resembles the effectiveness of RLHF (Ouyang et al., 2022) on top of supervised fine-tuning. (2) The binarized rewards () used in A3T lead to a significant improvement of success rate. We leave the incorporation of advanced RL (Zhou et al., 2024) and RLAIF (Lee et al., 2023; Chen et al., 2024; Hosseini et al., 2024; Yuan et al., 2024) algorithms into A3T as future work.
@@
While all of the experiments we previously reported are conducted with Mistral-7B-Instruct-v0.2 and QLoRA finetuning, in this section, we also validate A3T with gpt-3.5-turbo-1106 finetuning, the proprietary service provided by OpenAI. As the initial trajectory set for Round-0 training in A3T is obtained by ReAct prompting with gpt-3.5-turbo-instruct-0914, the starting point for the two base LLMs is the same.
Tables 8 and 9 report the performance comparison of Round-0 supervised training between the open-sourced and the proprietary LLMs. In AlfWorld, the performance of the QLoRA fine-tuned Mistral-7B-Instruct-v0.2 even surpasses that of the proprietary gpt-3.5-turbo-1106 fine-tuning service. In WebShop, the proprietary gpt-3.5-turbo-1106 finetuning performs better in Round-0 supervised training. We then let the two models separately compose diverse trajectories for their self-training. Because of the expense of inferring with the finetuned gpt-3.5-turbo-1106 model333The pricing is listed in https://openai.com/pricing, we compose trajectories for each failed training task ( with the open-sourced LLM). Shown in Table 10, the quality of the accumulated trajectories composed by the proprietary LLM is on par with those composed by the open-sourced LLM. After Round-1 self-training, the open-sourced model achieves an even higher test success rate. This is attributed to the proprietary service providing only the supervised fine-tuning option, while also indicating the importance of contrastive fine-tuning in A3T.